OpenStudy (anonymous):

A diffraction grating with 260 lines per mm is used in an experiment to study the visible spectrum of a gas discharge tube. At what angle from the beam axis will the first order peak occur if the tube emits light with wavelength of 580.2 nm? 8.676 deg Which I got to be =8.676degree

OpenStudy (anonymous):

But what would be part two..."At what angle will the second order peak occur?"

OpenStudy (joannablackwelder):

I agree with your answer. To find the second order peak, you need to make n=2.

OpenStudy (joannablackwelder):

\[\lambda=\frac{dsin \theta }{n }\]

OpenStudy (anonymous):

I did that and it says its wrong @JoannaBlackwelder

OpenStudy (joannablackwelder):

Hm, that's weird. Did you get 17.56 deg?

OpenStudy (anonymous):

no lol! How did you do that?? and thank you!!

OpenStudy (joannablackwelder):

Using the formula I gave you with n=2. :)

OpenStudy (anonymous):

huh...I must have been doing something wrong but thanks again! I really appreciate it :)

OpenStudy (joannablackwelder):

What did you get?

OpenStudy (joannablackwelder):

No worries. Happy studies!

OpenStudy (anonymous):

Okay so i am doing 2*580.2E-9/260 then taking the inverse sin of that but i cant get 17.60...could you show me the formula?

OpenStudy (joannablackwelder):

From the formula I gave you I plugged in λ = 580.2E-9, d=1/260,000 (to put it in slits/m), and n=2

OpenStudy (joannablackwelder):

Then solve as you suggested using inverse sin.

OpenStudy (anonymous):

Okay thank yoU!!!!